OpenStudy (anikate):

Find d if asub1=5 and a sub 60=300

OpenStudy (anikate):

\[a _{1}=5 & a _{60}=300\]

OpenStudy (anikate):

@karatechopper

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

what is d?

OpenStudy (anonymous):

Set up an equation using the nth term formula. \[a_{60} = a_1 + (n - 1)d\]\[300 = 5 + (60 - 1)d\]\[300 = 5 + 59d\]Can you solve for d (the common difference)?

OpenStudy (anikate):

ummm. just a min. Need to get calculator

OpenStudy (anikate):

5

OpenStudy (anikate):

umm. calc math

OpenStudy (anikate):

is it right?

OpenStudy (anonymous):

Yes.

OpenStudy (anikate):

oh goody. thx

OpenStudy (anonymous):

np :)